# FreeGraftor: Training-Free Cross-Image Feature Grafting for Subject-Driven Text-to-Image Generation

### v3

@@ 6. $I^{\text{gen}}\leftarrow\text{SAFG}(\epsilon^{\text{init}},\mathcal{F}^{\text{ref}},P)$ ;
Figure 4. Illustration of the Semantic-aware Feature Grafting (SAFG) Module. For each patch in the reference image, it first establishes semantic correspondence in the generated image via feature matching. The module then (1) concatenates the key and value pairs of the reference patch with those of the corresponding generated patch, while (2) copying and applying the position embedding from the generated patch to the reference patch’s key. This mechanism enables effective positional information sharing between corresponding patches while maintaining semantic alignment.
@@ 5. Compute revised attention:
3.2. Semantic-Aware Feature Grafting
Inspired by appearance transfer techniques(Epstein et al., 2023; Alaluf et al., 2024; Mou et al., 2023, 2024; Go et al., 2024), we leverage the inherent capabilities of text-to-image diffusion models to transfer visual characteristics of reference subjects to corresponding regions in generated images without incurring additional training costs. To achieve this, we propose the Semantic-Aware Feature Grafting (SAFG) module, which first establishes semantic correspondences between reference and generated patches via semantic matching, then fuses features through position-constrained attention fusion (Fig. 4).
3.2.1. Semantic Matching
Early appearance transfer methods(Alaluf et al., 2024; Mou et al., 2023, 2024) directly concatenate keys and values of reference and generated images to compute cross-image attention. However, high attention scores do not necessarily indicate semantic correspondences across images, and erroneous matches may lead to structural distortions and visual artifacts. DIFT(Tang et al., 2023) revealed that U-Net-based diffusion models contain rich semantic information in their feature space and established cross-image correspondences via cosine similarity-based semantic matching. We extend this principle to MM-DiT-based diffusion models.
Given a reference image $I^{\text{ref}}$ and a predefined binary mask $M^{\text{ref, pre}}$ indicating the reference subject, we extract attention features using rectified flow inversion(Rout et al., 2024; Wang et al., 2024b; Deng et al., 2024) and match reference $H^{\text{ref}}$ and generated $H^{\text{gen}}$ features at the same timestep and attention layer. For each reference patch $i$ within $M^{\text{ref, pre}}$, we compute cosine similarity with all generated patches $j$:
(4) $s_{ij}=\frac{H_{i}^{\text{img,ref}}\cdot H_{j}^{\text{img,gen}}}{\|H_{i}^{\text{img,ref}}\|_{2}\|H_{j}^{\text{img,gen}}\|_{2}},\quad m(i)=\arg\max_{j}s_{ij}$
(5) $M_{i}^{\text{ref, sim}}=\begin{cases}1,&s_{i,m(i)}\geq\tau\\ 0,&s_{i,m(i)}<\tau\end{cases}$
ii. Cycle Consistency Filtering. Following DreamMatcher(Nam et al., 2024), for each reference patch $i$ (satisfying $M_{i}^{\text{ref, pre}}=1$), we find the most similar reference patch $k$ to its matched generated patch $m(i)$:
(6) $m^{-1}(m(i))=\arg\max_{k:M_{k}^{\text{ref,pre}}=1}\frac{H_{m(i)}^{\text{img,gen}}\cdot H_{k}^{\text{img,ref}}}{\|H_{m(i)}^{\text{img,gen}}\|_{2}\|H_{k}^{\text{img,ref}}\|_{2}}$
(7) $d_{i}=\|\mathbf{p}_{i}^{\text{ref}}-\mathbf{p}_{m^{-1}(m(i))}^{\text{ref}}\|_{2},\quad M_{i}^{\text{ref, consi}}=\begin{cases}1,&d_{i}\leq\delta\\ 0,&d_{i}>\delta\end{cases}$
(8) $M^{\text{ref}}=M^{\text{ref, pre}}\odot M^{\text{ref, sim}}\odot M^{\text{ref, consi}}$
3.2.2. Position-Constrained Attention Fusion
(9) $K^{\text{cat}}=[K^{\text{txt}};K^{\text{img,gen}};K_{M}^{\text{img,ref}}],$
$V^{\text{cat}}=[V^{\text{txt}};V^{\text{img,gen}};V_{M}^{\text{img,ref}}]$
(10) $PE^{\text{cat}}=[PE^{\text{txt}};PE^{\text{img,gen}};\{PE_{m(i)}^{\text{img,gen}}\mid M_{i}^{\text{ref}}=1\}]$
The revised attention computation is formulated as:
(11) $\tilde{Q}=PE\odot Q,\quad\tilde{K}^{\text{cat}}=PE^{\text{cat}}\odot K^{\text{cat}}$
(12) $A^{+}=\text{softmax}\left(\frac{\tilde{Q}(\tilde{K}^{\text{cat}})^{\top}}{\sqrt{d}}\right),\quad H_{\text{out}}^{+}=A^{+}V^{\text{cat}}$
3.3. Structure-Consistent Initialization
To prevent the generated subject from overfitting the pose of the reference subject, we introduce a dynamic feature dropout strategy. Since structural layouts are predominantly determined during early diffusion steps(Balaji et al., 2022; Alaluf et al., 2024), we reduce the injection of reference features during the early steps while retaining them in the later steps to preserve visual details. At timestep $t$, for each reference patch, we apply dropout with probability $p=\omega t$, where $\omega\in[0,1]$ controls dropout intensity. The dropout mask is represented by a Bernoulli distribution:
(13) $M^{\text{drop}}_{i,j}\sim\text{Bernoulli}(1-\omega t)\quad\forall i,j$
The final reference mask is computed by applying dropout to the reference image mask obtained from semantic matching (Sec. 3.2):
(14) ${M^{\text{ref}}}^{\prime}=M^{\text{ref}}\odot M^{\text{drop}}$
This strategy preserves the reference subject’s structure while aligning it with the text-driven layout and maintaining pose flexibility. For multi-subject generation, we iteratively replace subjects in the template and invert the multi-subject collage. This avoids simultaneous processing of multiple reference images and eliminates additional computational overhead.
4. Experiments
4.1. Experimental Setup
4.1.1. Implementation Details
Our method is implemented on FLUX.1-dev. For collage construction, we use Grounding DINO(Liu et al., 2024) to localize subjects, SAM(Kirillov et al., 2023) to segment subjects, and LaMa(Suvorov et al., 2022) to erase the original subject in the template image. FireFlow(Deng et al., 2024) is employed for rectified flow inversion. Template generation, inversion, and final generation each take 25 diffusion steps. During semantic matching, we set the similarity threshold $\tau=0.2$, cycle consistency threshold $\delta=1.5$, and dropout intensity $\omega=0.5$. All experiments are conducted on an NVIDIA L20 GPU.
4.1.2. Baselines
We compare our method with zero-shot methods (BLIP-Diffusion, IP-Adapter(Ye et al., 2023), MS-Diffusion(Wang et al., 2024a), UNO(Wu et al., 2025), DreamO(Mou et al., 2025)) and training-free methods (FreeCustom(Ding et al., 2024), DiptychPrompting(Shin et al., 2024), LatentUnfold(Kang et al., 2025)). Detailed descriptions of these methods are provided below.
  • FreeCustom(Ding et al., 2024) provides a tuning-free approach for generating multi-concept images by leveraging a multi-reference self-attention mechanism along with a weighted mask strategy so that the generated image attends to each reference concept. With its weighted masking strategy to suppress extraneous details. This design efficiently preserves the semantic content of each input concept within context without further training.
@@ 4.1.2. Baselines
  • BLIP-Diffusion(Li et al., 2023) trains a multimodal encoder to produce a unified text-aligned visual representation of the target subject. During training, it learns to leverage this subject representation so the diffusion model can generate new renditions of the subject from text prompts.
  • MS-Diffusion(Wang et al., 2024a) injects reference images and explicit grounding tokens using a CLIP-based feature resampler to preserve fine details for each reference subject. Using the provided layout as guidance, it applies a multi-subject cross-attention mechanism so that each subject’s conditioning affects a specific region of the image, which harmonizes the composition of all subjects while respecting the text prompt.
@@
  • IP-Adapter(Ye et al., 2023) introduces a lightweight module that projects visual features from a CLIP image encoder directly into the cross-attention layers of a pre-trained text-to-image diffusion model. This enables the seamless integration of image prompts with textual conditions without the need for extra fine-tuning.
@@
  • UNO(Wu et al., 2025) first uses the inherent in-context generation ability of MM-DiTs to synthesize large-scale, high-consistency multi-subject paired data. Then it trains a “subject-to-image” model (conditioned on multiple images) with progressive cross-modal alignment and a universal rotary positional embedding. This pipeline yields a model that maintains high consistency while supporting control in both single- and multi-subject generation tasks.
@@
  • DreamO(Mou et al., 2025) is a unified image customization framework built on a MM-DiT. DreamO is trained on a large mixed dataset covering diverse tasks (identity, subject, style, background, etc.), and uses techniques to integrate multiple conditions. It employs a feature-routing constraint that queries the relevant information from each reference, and a placeholder strategy that assigns each condition a dedicated position in the input. DreamO can perform many customization tasks with high quality and flexibly combine different types of control conditions in one model.
@@
  • DiptychPrompting(Shin et al., 2024) reinterprets subject-driven generation as a text-conditioned inpainting task using a diptych layout, where the left panel holds a reference subject (with background removed) and the right panel is generated from a textual prompt. By enhancing the cross-attention between the two halves, DiptychPrompting precisely aligns the generated subject with the reference.
@@
  • LatentUnfold(Kang et al., 2025) reframes subject generation as grid-based image completion: the reference images are simply tiled in a mosaic layout to “activate” the model’s latent identity features. This simple “free lunch” approach is enhanced with a cascade attention design and a meta-prompting technique to boost fidelity.
4.1.3. Datasets and Evaluation
For qualitative evaluation, we use reference images from DreamBench(Ruiz et al., 2023), Mix-of-Show(Gu et al., 2023) and CustomConcept101(Kumari et al., 2023) for subject-driven generation. For quantitative evaluation, we generate 30×25×4=3,000 images using 30 subjects and 25 prompt templates from DreamBench(Ruiz et al., 2023), with 4 random seeds per subject-prompt pair. We evaluate performance from two perspectives: text alignment and image alignment. For text alignment, we report results using CLIP(Radford et al., 2021) and ImageReward(Xu et al., 2024) metrics. For image alignment, we use GroundedSAM(Ren et al., 2024) to extract segments of reference and generated subjects, respectively, and compute similarity scores between these segments using CLIP(Radford et al., 2021) and DINOv2(Oquab et al., 2023).
4.2. Quantitative Results
As shown in Table 1, methods using smaller base models (FreeCustom and BLIP-Diffusion) exhibit limitations in both image and text alignments. MS-Diffusion and IP-Adapter achieve lower image alignment than UNO and DreamO, suggesting that reliance solely on CLIP-based external encoders is insufficient for preserving intricate subject details. Training-free methods (DiptychPrompting and LatentUnfold) underperform in image alignment compared to trained zero-shot alternatives (UNO and DreamO), which indicates that the inherent context learning capacity of the base model itself becomes a limiting factor. Crucially, our method achieves the highest alignment across all metrics, demonstrating superior subject fidelity and textual adherence, validating its effectiveness for subject-driven generation.
Image-alignment Text-alignment Efficiency
Method Base Model CLIP-I DINO CLIP-T ImageReward Time (s) Memory (MiB)
FreeCustom SD v1.5 0.8308 0.6107 0.3246 0.6223 14.93 12068.73
BLIP-Diffusion SD v1.5 0.8621 0.7364 0.2949 0.2360 1.44 3641.84
MS-Diffusion SDXL 0.9023 0.7977 0.3254 1.3405 4.02 12268.41
IP-Adapter FLUX.1-dev 0.8920 0.7696 0.3048 0.7444 7.45 38690.51
UNO FLUX.1-dev 0.9069 0.8237 0.3151 1.2788 14.02 35248.43
DreamO FLUX.1-dev 0.9159 0.8400 0.3217 1.6234 12.15 34431.61
DiptychPrompting FLUX.1-dev 0.8924 0.7971 0.3291 1.5728 43.27 44928.55
LatentUnfold FLUX.1-dev 0.8741 0.7755 0.3170 1.3929 104.90 35402.22
Ours FLUX.1-dev $\mathbf{0.9527}$ $\mathbf{0.9042}$ $\mathbf{0.3308}$ $\mathbf{1.6481}$ 40.93 28993.33
Table 1. Quantitative Comparison of Subject-Driven Generation Methods. Without requiring model fine-tuning or additional training, our FreeGraftor significantly outperforms existing zero-shot and training-free approaches in both image-alignment and text-alignment.
In additon, we compare the efficiency of FreeGraftor with test-time subject-driven generation methods (including zero-shot and training-free approaches) on an NVIDIA L20 GPU. All methods generate 512×512 resolution images under identical conditions, with measurements of peak GPU memory consumption and inference time. As shown in Table 1, among methods built on the FLUX.1-dev base model, FreeGraftor significantly outperforms all other approaches in terms of memory usage, while also achieving shorter inference time compared to the other two training-free methods (DiptychPrompting and LatentUnfold). Compared to zero-shot methods, our approach is also competitive. While IP-Adapter, UNO, and DreamO have faster inference speeds than FreeGraftor, they consume more GPU memory. Due to their smaller parameter sizes, FreeCustom, BLIP Diffusion, and MS-Diffusion achieve high efficiency in both time and memory, but this comes at the cost of reduced subject consistency and lower visual quality. Notably, our FreeGraftor maintains an optimal balance between generation quality and computational efficiency, delivering robust performance without compromising subject faithfulness or visual details.
Figure 5. Generation results with single reference subject using different methods. Our FreeGraftor achieves pixel-level detail preservation (e.g., text and patterns) while allowing flexible text-guided control (e.g., poses of teddy bears and dogs).
Figure 6. Generated results with different text prompts under the same reference image.
Figure 7. Generation results with multiple reference subjects using different methods. Our FreeGraftor preserves the visual details of all reference subjects while producing high-quality images well-aligned with the text.
4.3. Qualitative Results
4.3.1. Single-Subject Generation
Fig. 5 presents the results for single-subject generation using FreeGraftor and other methods. FreeCustom, BLIP-Diffusion and IP-Adapter fail to retain fine visual details of the reference subject effectively. While other baselines partially recover reference characteristics, they struggle with fine-grained fidelity. FreeGraftor achieves pixel-level detail preservation, such as logos on the can, numbers on the clock, and text on the teddy bear. For non-rigid subjects (e.g., dogs, plush toys), FreeGraftor generates natural pose variations aligned with text guidance, demonstrating the flexibility and robustness of our feature grafting technique.
Extended single-subject generation results with various prompts are visualized in Fig. 6. Our method demonstrates comprehensive prompt adaptability, dynamically modulating subject attributes, poses, and scene configurations to synthesize high-fidelity images that achieve both visual harmony and precise semantic alignment with input textual descriptions.
4.3.2. Multi-Subject Generation
As shown in Fig. 7, FreeCustom and MS-Diffusion struggle to achieve sufficient subject consistency. While UNO and DreamO demonstrate higher fidelity, they suffer from subject omission (e.g., the barn and the table) and attribute confusion (e.g., confusion between the cat and the dog). FreeGraftor preserves all reference details while enabling flexible text-guided control, highlighting its effectiveness in multi-subject generation.
Figure 8. Generation results of different ablation variants. (a) Without structure-consistent initialization. (b) Without feature grafting. (c) Using direct key-value concatenation in attention layers instead of semantic matching. (d) Using direct patch feature replacement instead of position-constrained attention fusion. (e) Our full method.
Image-alignment Text-alignment
Variant CLIP-I DINO CLIP-T ImageReward
w/o Structure-consistent Initialization 0.8529 0.7138 0.3236 1.6320
w/o Feature Grafting 0.8463 0.6572 0.3306 1.6434
w/o Semantic Matching 0.9504 0.8974 0.3232 1.5258
w/ Feature Replacement 0.8476 0.7010 0.3174 1.4551
Ours $\mathbf{0.9527}$ $\mathbf{0.9042}$ $\mathbf{0.3308}$ $\mathbf{1.6481}$
Table 2. Quantitative Comparison of Ablation Variants. Absence of structure-consistent initialization or feature grafting leads to degraded image alignment, while omitting semantic matching or replacing patch features results in deteriorated text alignment.
4.4. Ablation Study
4.4.1. Component Ablation
To validate the effectiveness of each component in our method, we compare several ablation variants, as shown in Fig. 8. (a) When structure-consistent initialization is removed, the generated images inherit the reference subject’s appearance (e.g., colors and patterns) but fail to maintain structural consistency (e.g., distorted body proportions in dogs). (b) Without feature grafting, there is no visual correlation between the generated and reference subjects. (c) If semantic matching is omitted and reference keys/values are directly appended during attention computation, the generated subjects exhibit rigid poses and artifacts. (d) Replacing position-constrained attention fusion with direct feature replacement introduces incoherent edges and distortions. Quantitative results in Table 2 confirm that both structure-consistent initialization and feature grafting significantly improve image alignment, as both are critical for preserving subject identity. Skipping semantic matching or positional fusion degrades text alignment due to reduced image quality.
Figure 9. Generation results under different similarity threshold ($\tau$) and cycle-consistency threshold ($\delta$) settings. Higher $\tau$ and lower $\delta$ lead to more stringent filtering.:
Figure 10. Generation results under different dropout intensity ($\omega$) settings. Higher $\omega$ improves text alignment but degrades subject consistency.:
4.4.2. Hyperparameter Analysis
@@
We conduct additional ablation experiments to analyze the impact of similarity threshold $\tau$ and cycle consistency threshold $\delta$ on semantic matching. Appropriate thresholds alleviate mismatches (e.g., unnatural expansion of the backpack), while overly strict thresholds ($\tau\uparrow$ or $\delta\downarrow$) discard critical visual details (e.g., missing badges on the backpack, Fig. 9).
We further verify the effectiveness of dropout in mitigating structural overfitting. As shown in Fig. 10, moderate dropout enhances pose flexibility while preserving subject identity, whereas excessive dropout intensities ($\omega>0.6$) erode identity retention. These findings highlight the necessity of balancing pose flexibility and detail preservation through parameter tuning.
Figure 11. Comparison between DreamBooth and our method on FLUX.1. Without computation-intensive fine-tuning, our method achieves performance comparable to DreamBooth and even surpasses it in capturing visual details such as text and patterns.
Figure 12. Generation results of (a) Stable Diffusion (b)Stable Diffusion with DreamMatcher (c) DreamBooth on Stable Diffusion (d) DreamBooth on Stable Diffusion with DreamMatcher (e) Ours.
4.5. Comparison with DreamBooth
@@
We compare FreeGraftor with DreamBooth, a widely-used tuning-based method. Due to limitations in computational resources, we implemented DreamBooth using LoRA instead of full parameter fine-tuning on FLUX.1-Dev, with a LoRA rank of 16.
@@
As shown in Fig. 11, although fine-tuned for several tens of minutes on the reference images of each subject, DreamBooth still struggles to fully reproduce the visual details of the reference subject (e.g., text and patterns). Additionally, fine-tuning on limited samples can lead to overfitting of the background, compromising its ability to follow textual semantics. In contrast, without resource-intensive fine-tuning, our method achieves pixel-level detail preservation and faithful text alignment.
Image-alignment Text-alignment
Variant CLIP-I DINO CLIP-T ImageReward
DreamBooth 0.9261 0.8776 0.3123 1.3232
Ours $\mathbf{0.9527}$ $\mathbf{0.9042}$ $\mathbf{0.3308}$ $\mathbf{1.6481}$
Table 3. Quantitative Comparison of DreamBooth and our method on FLUX.1-Dev.
Furthermore, we quantitatively compared FreeGraftor with DreamBooth using the same experimental setup (Sec. 4.1.3). Table 3 shows our approach delivers significant performance advantages beyond just efficiency gains. FreeGraftor resolves the fidelity-efficiency dilemma in subject-driven generation, going beyond merely alleviating the limitations of training-free methods.
4.6. Comparison with DreamMatcher
@@
The most comparable approach to ours is likely DreamMatcher(Nam et al., 2024), yet the two remain fundamentally distinct. Firstly, DreamMatcher is not a standalone subject-driven generation method but rather a plugin. It requires integration with tuning-based approaches (e.g., DreamBooth), which involve time-consuming and resource-intensive test-time optimization. In contrast, our method is a standalone subject-driven generation pipeline without any fine-tuning or training. Secondly, DreamMatcher adopts a direct attention feature replacement mechanism for identity injection, which tends to introduce artifacts along subject boundaries. In contrast, our approach leverages cross-image attention with shared spatial information, employing a more refined methodology that effectively mitigates image distortion.
@@
Image-alignment Text-alignment
Variant CLIP-I DINO CLIP-T ImageReward
Stable Diffusion + DreamMatcher 0.7830 0.5336 0.3232 0.6387
DreamBooth + DreamMatcher 0.8912 0.7918 0.3021 0.1767
Ours $\mathbf{0.9527}$ $\mathbf{0.9042}$ $\mathbf{0.3308}$ $\mathbf{1.6481}$
Table 4. Quantitative Comparison of DreamMatcher and our method.
Quantitative comparisons under identical settings (Sec. 4.1.3) confirm our superiority, as shown in Table 4. Our method achieves significant performance gains through two key designs. First, the structure-consistent noise initialization produces reasonable image layouts while faithfully preserving reference geometry, which ensures robust semantic matching and detail retention. Second, our position-constrained attention fusion mechanism enables harmonious integration of features, resulting in more natural-looking outputs. Ablation studies (Sec. 4.4.1) confirm the effectiveness of these core components.
5. Conclusion
We present FreeGraftor, a novel training-free framework for subject-driven generation that bridges the critical gap between subject fidelity and computational efficiency. By leveraging Semantic-Aware Feature Grafting, FreeGraftor seamlessly transfers the appearance of reference subjects into generated scenes. Additionally, our Structure-Consistent Initialization strategy preserves reference geometry and enhances the robustness of semantic matching. Extensive experiments demonstrate that FreeGraftor achieves pixel-level detail preservation and flexible text-guided control while requiring no training or test-time optimization. Our framework can be seamlessly extended to multi-subject generation, faithfully retaining visual details of all reference subjects while incurring no additional computational overhead.
Acknowledgements.
This work was supported by Beijing Natural Science Foundation (L247010), NSFC (No.61906018).

### v4

@@ 6. $I^{\text{gen}}\leftarrow\text{SAFG}(\epsilon^{\text{init}},\mathcal{F}^{\text{ref}},P)$ ;
.0-A2 Semantic-Aware Feature Grafting
Algorithm 2 details the Semantic-Aware Feature Grafting (SAFG) module, including semantic matching and position-constrained attention fusion.
@@ 5. Compute revised attention:
.0-B Baseline Methods
.0-B1 Zero-Shot Methods
Zero-shot methods do not require any subject-specific training or fine-tuning, making them efficient but often less effective at preserving subject details.
@@
  • BLIP-Diffusion[li2023blip] trains a multimodal encoder to produce a unified text-aligned visual representation of the target subject. During training, it learns to leverage this subject representation so the diffusion model can generate new renditions of the subject from text prompts.
@@
  • IP-Adapter[ye2023ip] introduces a lightweight module that projects visual features from a CLIP image encoder directly into the cross-attention layers of a pre-trained text-to-image diffusion model. This enables the seamless integration of image prompts with textual conditions without the need for extra fine-tuning.
  • MS-Diffusion[wang2024ms] injects reference images and explicit grounding tokens using a CLIP-based feature resampler to preserve fine details for each reference subject. Using the provided layout as guidance, it applies a multi-subject cross-attention mechanism so that each subject’s conditioning affects a specific region of the image, which harmonizes the composition of all subjects while respecting the text prompt.
@@
  • UNO[wu2025less] first uses the inherent in-context generation ability of MM-DiTs to synthesize large-scale, high-consistency multi-subject paired data. Then it trains a ”subject-to-image” model (conditioned on multiple images) with progressive cross-modal alignment and a universal rotary positional embedding. This pipeline yields a model that maintains high consistency while supporting control in both single- and multi-subject generation tasks.
@@
  • DreamO[mou2025dreamo] is a unified image customization framework built on a MM-DiT. DreamO is trained on a large mixed dataset covering diverse tasks (identity, subject, style, background, etc.), and uses techniques to integrate multiple conditions. It employs a feature-routing constraint that queries the relevant information from each reference, and a placeholder strategy that assigns each condition a dedicated position in the input. DreamO can perform many customization tasks with high quality and flexibly combine different types of control conditions in one model.
.0-B2 Training-Free Methods
Training-free methods avoid any form of training or fine-tuning, operating directly on pre-trained models.
  • FreeCustom[ding2024freecustom] provides a tuning-free approach for generating multi-concept images by leveraging a multi-reference self-attention mechanism along with a weighted mask strategy so that the generated image attends to each reference concept. With its weighted masking strategy to suppress extraneous details, this design efficiently preserves the semantic content of each input concept within context without further training.
@@
  • DiptychPrompting[shin2024large] reinterprets subject-driven generation as a text-conditioned inpainting task using a diptych layout, where the left panel holds a reference subject (with background removed) and the right panel is generated from a textual prompt. By enhancing the cross-attention between the two halves, DiptychPrompting precisely aligns the generated subject with the reference.
@@
  • LatentUnfold[kang2025flux] reframes subject generation as grid-based image completion: the reference images are simply tiled in a mosaic layout to ”activate” the model’s latent identity features. This simple ”free lunch” approach is enhanced with a cascade attention design and a meta-prompting technique to boost fidelity.
.0-C Additional Experiments
.0-C1 Hyperparameter Analysis
We analyze the impact of key hyperparameters on performance, including consistency thresholds and dropout intensity.
@@
Fig. 6 shows generation results under different similarity threshold ($\tau$) and cycle consistency threshold ($\delta$) settings. Appropriate thresholds alleviate mismatches (e.g., unnatural expansion of the backpack), while overly strict thresholds ($\tau\uparrow$ or $\delta\downarrow$) discard critical visual details (e.g., missing badges on the backpack).
Figure 6: Generation results under different similarity threshold ($\tau$) and cycle-consistency threshold ($\delta$) settings. Higher $\tau$ and lower $\delta$ lead to more stringent filtering.
Fig. 7 shows the effect of dropout intensity ($\omega$). Moderate dropout ($\omega=0.5$) enhances pose flexibility while preserving subject identity. Excessive dropout ($\omega>0.6$) erodes identity retention, while insufficient dropout limits pose variation.
Figure 7: Generation results under different dropout intensity ($\omega$) settings. Moderate dropout balances pose flexibility and detail preservation.
.0-C2 Multi-Subject Generation Results
FreeGraftor can seamlessly extend to multi-subject generation. As shown in Fig. 8, FreeCustom and MS-Diffusion struggle to achieve sufficient subject consistency. While UNO and DreamO demonstrate higher fidelity, they suffer from subject omission (e.g., the barn and the table) and attribute confusion (e.g., confusion between the cat and the dog). FreeGraftor preserves visual details of all reference subjects while producing high-quality images well-aligned with the text, highlighting its effectiveness in multi-subject generation.
Figure 8: Multi-subject generation results. FreeGraftor preserves all reference details without layout conflicts or attribute confusion.
.0-C3 Efficiency Analysis
Table III compares the efficiency of FreeGraftor with other methods on an NVIDIA L20 GPU. All methods generate 512×512 resolution images under identical conditions.
TABLE III: Efficiency Comparison of Subject-Driven Generation Methods
Method Base Model Time (s) Memory (MiB)
FreeCustom Stable Diffusion v1.5 14.93 12068.73
BLIP-Diffusion Stable Diffusion v1.5 1.44 3641.84
MS-Diffusion SDXL 4.02 12268.41
IP-Adapter FLUX.1-dev 7.45 38690.51
UNO FLUX.1-dev 14.02 35248.43
DreamO FLUX.1-dev 12.15 34431.61
DiptychPrompting FLUX.1-dev 43.27 44928.55
LatentUnfold FLUX.1-dev 104.90 35402.22
Ours FLUX.1-dev 40.93 28993.33
As shown in Table III, among methods built on the FLUX.1-dev base model, FreeGraftor significantly outperforms all other approaches in terms of memory usage, while also achieving shorter inference time compared to the other two training-free methods (DiptychPrompting and LatentUnfold). Compared to zero-shot methods, our approach is also competitive. While IP-Adapter, UNO, and DreamO have faster inference speeds than FreeGraftor, they consume more GPU memory. Due to their smaller parameter sizes, FreeCustom, BLIP Diffusion, and MS-Diffusion achieve high efficiency in both time and memory, but this comes at the cost of reduced subject consistency and lower visual quality. Notably, our FreeGraftor maintains an optimal balance between generation quality and computational efficiency, delivering robust performance without compromising subject faithfulness or visual details.
.0-C4 Comparison with DreamBooth
@@
We compare FreeGraftor with DreamBooth, a widely-used tuning-based method. We implemented DreamBooth using LoRA (rank 16) on FLUX.1-Dev due to computational constraints.
@@
Fig. 9 shows qualitative results. Although fine-tuned for tens of minutes on each subject’s reference images, DreamBooth struggles to fully reproduce visual details (e.g., text and patterns) and can overfit to background elements, compromising text alignment. In contrast, FreeGraftor achieves pixel-level detail preservation and faithful text alignment without resource-intensive fine-tuning.
Figure 9: Comparison between DreamBooth and our method on FLUX.1. Without computation-intensive fine-tuning, our method achieves superior detail preservation.
.0-C5 Comparison with DreamMatcher
@@
The most comparable approach to ours is likely DreamMatcher[nam2024dreammatcher], yet the two remain fundamentally distinct. Firstly, DreamMatcher is not a standalone subject-driven generation method but rather a plugin. It requires integration with tuning-based approaches (e.g., DreamBooth), which involve time-consuming and resource-intensive test-time optimization. In contrast, our method is a standalone subject-driven generation pipeline without any fine-tuning or training. Secondly, DreamMatcher adopts a direct attention feature replacement mechanism for identity injection, which tends to introduce artifacts along subject boundaries. In contrast, our approach leverages cross-image attention with shared spatial information, employing a more refined methodology that effectively mitigates image distortion.
@@
Figure 10: Generation results of (a) Stable Diffusion (b) Stable Diffusion with DreamMatcher (c) DreamBooth on Stable Diffusion (d) DreamBooth on Stable Diffusion with DreamMatcher (e) Ours.
.0-C6 Prompt Adaptability Results
Fig. 11 demonstrates FreeGraftor’s comprehensive prompt adaptability. Our method dynamically modulates subject attributes, poses, and scene configurations while maintaining pixel-level detail preservation, achieving both visual harmony and precise semantic alignment with input textual descriptions.
Figure 11: Extended single-subject generation results with various prompts. FreeGraftor demonstrates robust prompt adaptability while preserving subject identity.
.0-D Social Impact
The proposed FreeGraftor framework has significant potential to democratize personalized content creation by enabling users without professional technical backgrounds to synthesize high-fidelity images of specific subjects without extensive training. This low-threshold advantage eliminates costly computing resource barriers that long excluded non-experts, empowering creators in digital art, advertising, and education to rapidly prototype visuals—such as custom art designs, brand-aligned product images, and teaching illustrations—while preserving intricate details, thereby lowering access to professional-grade content generation.
However, the ease of transferring visual identities raises prominent misuse concerns, including the generation of deceptive content like deepfakes for disinformation campaigns and unauthorized replication of copyrighted works that threaten intellectual property rights. Beyond ethical risks, reliance on large pre-trained diffusion models still causes significant energy consumption during inference—especially for high-resolution images—highlighting the need for efficient deployment research such as model compression.
To address these challenges, we advocate for ethical guidelines balancing creative freedom and anti-malicious safeguards. Specific measures include embedding imperceptible watermarks for traceability, implementing access controls for sensitive reference data like private portraits, and fostering industry collaboration to establish responsible use standards.